# Three-dimensional Abelian and nonAbelian gauge Higgs theories

### v1

@@ A.3.2 Finite-size scaling of the order-parameter correlation function
where $y_{\phi}$ is the RG dimension of the field $\phi({\bm{x}})$, ${\bm{x}}\equiv{\bm{x}}_{2}-{\bm{x}}_{2}$ and we assumed translation invariance (i.e., the absence the boundaries). The integral of $G({\bm{x}},L,r)$ over the whole finite volume $L^{d}$ provides the magnetic susceptibility
@@
A.3.3 Finite-size scaling of renormalization-group invariant quantities
Dimensionless RG invariant quantities are particularly useful to investigate the critical region. Examples of such quantities are the ratio
@@ A.3.3 Finite-size scaling of renormalization-group invariant quantities
(or the analogous ones with $<$ instead of $>$), one can define $r_{\rm cross}$ by requiring
@@
The crossing point $r_{\rm cross}$ converges to $r=0$, with corrections that are typically suppressed as $L^{-1/\nu-\omega}$.
We finally note that using a RG invariant quantity $\widetilde{R}$ that is monotonic with respect to the relevant parameter $r$, as is generally the case of $R_{\xi}=\xi/L$, one may write a second generic RG invariant quantity $R_{i}$ as
$R_{i}(L,r)=\widehat{\cal R}_{i}(R_{\xi})+O(L^{-\omega}),$ (152)
where ^$\widehat{\cal R}_{i}$ depends on the universality class only, without any nonuniversal normalization of the argument. This is true once the boundary conditions and the shape of the lattice have been fixed. The scaling relation (152) is particularly convenient to test universality-class predictions, since it permits easy comparisons between different models without the need of any tuning of nonuniversal parameters. One may write the FSS behavior of any other quantity in terms of the $R_{\xi}$. For example, the asymptotic FSS behavior of the susceptibility $\chi$ defined in Eq. (144) can be rewritten as
@@
A.3.4 Finite-size scaling of the energy cumulants
As discussed in this review, statistical models with gauge symmetries also develop topological transitions characterized by the absence of a local order parameter. We mention the continuous transitions belonging the the 3D ${\mathbb{Z}}_{Q}$ gauge and IXY universality classes that emerge in the LAH models, see Secs. 5 and 6. The numerical study of the critical properties at these topological transitions can be performed by analyzing the scaling properties of the energy cumulants, which are always gauge invariant and well defined, see, e.g., Refs. [63, 86, 174, 92].
where $Z$ is the partition function defined in Eq. (133). [^52] The cumulants $C_{k}$ can be related to the energy central momenta $M_{k}=\langle\,(H-\langle H\rangle)^{k}\,\rangle$, by $C_{1}=-\langle H\rangle/V$, $C_{2}=M_{2}/V$, $C_{3}=-L^{-3}M_{3}$, $C_{4}=(M_{4}-3M_{2}^{2})/V$, etc. Note that $C_{1}=-E$ where $E=\langle H\rangle/V$ is the energy density, and $C_{2}$ is proportional to the specific heat.
As one can easily check, the first cumulant $C_{1}=-E$ is dominated by the analytic background contribution $B_{1}$, due to the relation $\nu>1/d>0$, which is generally satisfied at continuous transitions [95]. This makes the scaling behavior of the energy density subleading, and therefore rather difficult to observe (this issue has been addressed in Refs. [499, 500]). Also the second cumulant $C_{2}$, which is proportional to the specific heat, is dominated by the analytic background when the specific-heat exponent $\alpha=2\nu-d$ is negative. The singular scaling part is always the dominant term in the higher energy cumulants $C_{k}$, for $k\geq 3$. Taking into account that their numerical computation becomes harder and harder with increasing $k$, due to significant cancellations which make their relative error quite large and increasing with $k$, optimal results are obtained by focusing on $C_{3}$ and $C_{4}$.
@@ A.4 Scaling behavior at a multicritical point
We now briefly outline of the FSS behavior at a multicritical point (MCP), see Figs. 1, which is generally characterized by the presence of two even relevant RG perturbations, which must be both tuned to approach the MCP (we assume that $r_{1}=r_{2}=0$ at the MCP). At a MCP, the singular part of the free-energy density can be written as [170, 171, 172, 95, 173, 174]
@@
A.5 $T\to 0$ scaling behavior of two-dimensional models with continuous symmetry
@@
These statistical systems are asymptotically free with a nonperturbatively generated mass gap. Such a property is also present in QCD, i.e. the theory of strong interactions, and thus these 2D theories have been often used as toy models to understand some of the nonperturbative mechanisms characterizing QCD, see, e.g., Refs. [436, 2, 95, 439]. We mention that the 2D $N$-vector model is also important in condensed-matter physics. For $N=3$ it describes the behavior for $T\to 0$ of the 2D spin-$S$ Heisenberg quantum antiferromagnet [429]. Indeed, at finite temperature $T$ this quantum spin system is described by a (2+1)-dimensional O(3) classical theory in which the Euclidean time direction has a finite extent $1/T$. In the critical limit the relation $1/T\ll\xi$ is satisfied, therefore the system becomes effectively two-dimensional, and thus its critical behavior is described by the 2D three-vector model.
The long-distance critical behavior in the zero-temperature limit can be predicted by performing perturbative computations in powers of $T$, see, e.g., Refs. [421, 422, 423, 424, 425, 426, 280, 427, 428, 429, 430, 431, 432, 433, 434, 435, 436, 2, 95]. In particular, the $T$-dependence of the correlation length $\xi$ can be inferred from the perturbative expansion of the lattice $\beta$ function, related with the derivative of the correlation length $\xi$ with respect to $T$,
@@
where $\xi_{0}$ is a nonperturbative model-dependent constant (depending also on the actual definition of $\xi$). More specifically, using the first known terms of the $\beta$ functions, the asymptotic behavior of the length scale in the zero-temperature limit is (see e.g. Refs. [2, 95])
@@
where $p$ is a model-dependent power. Therefore, apart from logarithmic corrections (due to $T\sim 1/\ln\xi$), the exponent associated with the leading scaling corrections is given by $\omega=2$. Such a behavior has been verified explicitly in the large-$N$ limit, see, e.g., Ref. [501].
We also mention that the nature of the asymptotic low-temperature behavior of the 2D RPN-1 models has been long debated, see e.g., Refs. [221, 219, 218, 217, 216, 215]. Refs. [217, 218, 219] reported arguments that support the claim that RPN-1 and $N$-vector models belong to the same universality class, implying the irrelevance of the ${\mathbb{Z}}_{2}$ gauge symmetry. They are essentially related to the fact that $N$-vector and RPN-1 have the same asymptotic perturbative behavior. However, extensive numerical results for several variant RPN-1 models [221], show substantial differences in the nonperturbative scaling functions, which may be explained by the nonperturbative relevance of the topological ${\mathbb{Z}}_{2}$ defects characterizing the RPN-1 models, giving rise to distinct universality classes.
We finally mention that 2D systems with an Abelian O(2) global symmetry, such as the $N=2$ XY vector model and the RP1 model, are peculiar in this respect, since they can undergo a finite-temperature topological BKT transition [413, 414, 415], which separates the high-$T$ disordered phase from the low-temperature nonmagnetized phase characterized by spin waves and quasi-long range order with vanishing magnetization. At BKT transitions the correlation length $\xi$ scales as $\xi\sim\exp(c/\sqrt{T-T_{c}})$ approaching the BKT critical temperature $T_{c}$ from the high-temperature phase. The scaling behavior at BKT transitions has been largely discussed in the literature, see, e.g., Refs. [413, 414, 415, 440, 441, 442, 443, 95, 444, 169].
@@ A.6 Finite-size scaling at first-order transitions
The behaviors emerging at first-order phase transitions are more complex than those observed at continuous transitions, see, e.g., Refs.[115, 117]. Critical phenomena at continuous phase transitions are essentially related with the presence of critical correlations, which decay as a power of the distance at the critical point, and with a diverging length scale $\xi$. When approaching the critical point, the long-distance behavior, i.e., on distances of the order of the scale $\xi$, shows universal features that only depend on few global properties of the microscopic short-distance interactions. The behavior at first-order transitions is instead more complex. Indeed, we typically observe the coexistence of different phases, with one or more phases that are ordered on all length scales, apart from some short-range local fluctuations. The presence of coexisting phases gives rise to peculiar competing phenomena in the phase-coexistence region, such as metastability, nucleation, droplet formation, coarsening, etc. [115, 502] Moreover, the observed bulk behavior crucially depends on the nature of the boundary conditions, at variance with what happens at continuous transitions, where boundary conditions only affect some finite-size properties of the system, but are irrelevant in the thermodynamic limit. The sensitivity to the boundary conditions represents one of the main qualitative differences between continuous and first-order phase transitions [117].
@@
To distinguish continuous and first-order transitions in numerical studies of systems with limited size, one can exploit the substantial differences of their FSS behavior. A possible approach consists in considering the large-$L$ behavior of the specific heat or of the susceptibility of the order parameter, that should have a maximum that diverges as $L^{d}$[511, 512, 513, 514] In practice, one should compute the maximum of the susceptibility $\chi_{A,\rm max}(L)$ of a local operator $A_{\bm{x}}$ for each value of $L$ and then fit the results to a power law $aL^{p}$. If one obtains $p\approx d$, the transition is of first order; otherwise, the transition is continuous, $p$ being related to standard critical exponents. If $A$ is the energy density, then $\chi_{A}$ is the specific heat that diverges as $L^{\alpha/\nu}$ (if $\alpha$ is positive), while if $A$ is the magnetization, then $\chi_{A}\sim L^{2-\eta}$. This approach works nicely for strongly discontinuous transitions. In the case of weakly discontinuous transition, however, the asymptotic behavior $\chi_{A}\sim L^{d}$ may set in for values of $L$ that are much larger than those considered in the simulations. Thus, data may show an effective scaling $\chi_{A,\rm max}(L)\sim L^{p}$, with $p$ significantly smaller than $d$, effectively mimicking a continuous transition (the Potts model [515, 516] in two and three dimensions is a good example, see Refs. [511, 512, 513, 514]).
@@
An extended review of FSS behaviors at thermal and quantum first-order transitions, and their dependence on the boundary conditions, can be found in Ref. [117].
@@ Appendix B Critical exponents of some O($N$)-vector universality classes
In this section we report the presently most accurate estimates of the critical exponents of the 3D $N$-vector universality classes, focusing on the most commen cases, the 3D Ising, XY, and O(3) vector universality classes. These universality classes are associated with the $\Phi^{4}$ field theory for an $N$-component real field. A a simple lattice representative is the standard $N$-vector model defined by the Hamiltonian (1).
Some accurate estimates of the critical exponents for the 3D Ising ($N=1$), XY ($N=2$), and Heisenberg ($N=3$) universality classes are reported in Tables 1, 2, and 3, respectively. A more complete list of theoretical and experimental results can be found in Ref. [95] (at least up to 2002); those reported here are the most accurate in the various approaches. We recall that the critical exponents controlling the asymptotic exponents of other observables can be obtained from $\nu$ and $\eta$ by using scaling and hyperscaling relations (see, e.g., Refs. [2, 114]). Accurate estimates of the critical exponents of the O($N$) vector universality classes for $N\geq 4$ can be found in Refs. [294, 95].
@@
Table 1: Estimates of the critical exponents of the 3D Ising universality class (characterized by a global ${\mathbb{Z}}_{2}$ symmetry). We report the correlation-length exponent $\nu$, the order-parameter exponent $\eta$, and the exponent $\omega$ associated with the leading scaling corrections. We report: SFT results obtained by resumming high-order perturbative expansions (6-loop calculations in the fixed-dimension scheme [517, 523, 524, 525, 526, 527] and in the $\varepsilon$-expansion scheme [518, 528, 529, 530]), by using nonperturbative RG approaches [519], and the conformal field theory (CFT) bootstrap [252, 531] approach; results obtained in lattice models by resumming high-temperature expansions [520] and in MC simulations [522, 521, 532, 533]. We note that there is an overall agreement among the results obtained by the different approaches. Moreover, there is also a good agreement with experiments in physical systems that undergo continuous transitions in the Ising universality class (liquid-vapor systems, binary systems, uniaxial magnetic systems, Coulombic systems, etc.) see, e.g., Ref. [95] for a list of experimental results.
@@
Table 2: Estimates of the critical exponents of the 3D XY universality class (characterized by a global O(2) symmetry). We report the correlation-length exponent $\nu$, the order-parameter exponent $\eta$, and the exponent $\omega$ associated with the leading scaling corrections. We report: SFT results obtained by resumming high-order perturbative expansions [517, 518] (6-loop calculations in the fixed-dimension and in the $\varepsilon$-expansion scheme), by using nonperturbative RG approaches [519] and the conformal field theory (CFT) bootstrap [321] approach; results obtained in lattice models by combining high-temperature and MC results [289] and in MC simulations [534, 288]. We also mention that experimental estimates in a microgravity environment have been obtained for the 4He superfluid transition [535, 536, 537]. See Ref. [95] for a more complete list of theoretical and experimental results.
@@
Table 3: Estimates of the critical exponents of the 3D Heisenberg universality class (characterized by a global O(3) symmetry). We report the correlation-length exponent $\nu$, the order-parameter exponent $\eta$, and the exponent $\omega$ associated with the leading scaling corrections. We report: SFT results obtained by resumming high-order perturbative expansions [517, 518] (6-loop calculations in the fixed-dimension and in the $\varepsilon$-expansion scheme), by using nonperturbative RG approaches [519], and the conformal field theory (CFT) bootstrap [538] approach; results obtained in lattice models by combining high-temperature and MC results [539, 250] and in MC simulations [540]. See Refs. [95, 540] for a more complete list of theoretical and experimental results.
Appendix C Mean-field analyses of the low-temperature Higgs phases
In this appendix we present a mean-field analysis of the low-temperature Higgs phases of the LNAH model with SU($N_{c}$) gauge symmetry and $N_{f}$ degenerate scalar flavors in the fundamental representation. This model has been discussed in Sec. 9.1. For more details see Ref. [99]. This analysis allows one to infer the symmetry properties of the low-temperature Higgs phases, which depend on the parameter $v$, the number of colors $N_{c}$ and of flavors $N_{f}$. It is worth noting that this discussion applies to generic $d$-dimensional systems, therefore also to $d=4$ space-time systems that may be relevant in the context of high-energy physics. Analogous analyses can be performed for SU($N_{c}$) theories with scalar matter in the adjoint representation [96, 259, 98], and SO($N_{c}$) gauge LNAH models [376].

### v3

@@ A.3.3 The order-parameter correlation function
where $y_{\phi}$ is the RG dimension of the field $\phi({\bm{x}})$, ${\bm{x}}\equiv{\bm{x}}_{1}-{\bm{x}}_{2}$, and we assumed translation invariance (i.e., the absence the boundaries). The integral of $G({\bm{x}},L,r)$ over the whole finite volume $L^{d}$ provides the magnetic susceptibility
@@
A.3.4 Renormalization-group invariant quantities
Dimensionless RG invariant quantities are particularly useful for the investigation of the critical behavior. Examples of such quantities are the ratio
@@ A.3.4 Renormalization-group invariant quantities
(or the analogous ones with $<$ replacing $>$), one can define $r_{\rm cross}$ by requiring
@@
The crossing point $r_{\rm cross}$ converges to $r=0$, with corrections that typically vanish as $L^{-1/\nu-\omega}$ as $L\to\infty$.
We finally note that, if a RG invariant quantity $R_{m}$ is monotonic with respect to the relevant parameter $r$, as is generally the case for $R_{\xi}=\xi/L$, one may write for a second generic RG invariant quantity $R_{i}$ the scaling relation
$R_{i}(L,r)=\widehat{\cal R}_{i}(R_{m})+O(L^{-\omega}),$ (157)
where ^$\widehat{\cal R}_{i}$ depends on the universality class only, without any nonuniversal normalization of the argument. This is true once the boundary conditions and the shape of the lattice have been fixed. The scaling relation (157) is particularly convenient to test universality-class predictions, since it allows one to compare the results for different models without the need of tuning nonuniversal parameters. One may write the FSS behavior of any other quantity in terms of $R_{\xi}$. For example, the asymptotic FSS behavior of the susceptibility $\chi$ defined in Eq. (149) can be rewritten as
@@ A.4 Scaling behavior at a multicritical point
We now briefly outline the FSS behavior at a multicritical point (MCP), see Fig. 1, which is generally characterized by the presence of two relevant RG perturbations, which are both invariant under the global symmetry of the model and which must be both tuned to approach the MCP. We indicate with $r_{1}$ and $r_{2}$ the corresponding Hamiltonian parameters, normalized so that $r_{1}=r_{2}=0$ at the MCP. At a MCP, the singular part of the free-energy density can be written as [201, 202, 203, 96, 204, 205]
@@
A.5 Two-dimensional spin models with continuous symmetry
@@
These statistical systems are asymptotically free with a nonperturbatively generated mass gap. Such a property is also present in QCD, the theory of strong interactions, and thus these 2D theories have often been used as toy models to understand some of the nonperturbative mechanisms characterizing QCD, see, e.g., Refs. [480, 2, 96, 483]. We mention that the 2D $N$-vector model is also important in condensed-matter physics. For $N=3$ it describes the zero-temperature behavior of the 2D spin-$S$ Heisenberg quantum antiferromagnet [473]. Indeed, at finite temperature $T$ this quantum spin system is described by a (2+1)-dimensional O(3) classical theory in which the Euclidean time direction has a finite extent $1/T$. In the critical limit the relation $1/T\ll\xi$ is satisfied, therefore the system becomes effectively two-dimensional, and thus its critical behavior is described by the 2D O(3)-vector model.
The long-distance critical behavior in the zero-temperature limit can be predicted by performing perturbative computations in powers of $T$, see, e.g., Refs. [465, 466, 467, 468, 469, 470, 311, 471, 472, 473, 474, 475, 476, 477, 478, 479, 480, 2, 96]. In particular, the $T$-dependence of the correlation length $\xi$ can be inferred from the perturbative expansion of the lattice $\beta$ function, defined in terms of the derivative of the correlation length $\xi$ with respect to $T$,
@@
where $\xi_{0}$ is a nonperturbative model-dependent constant (depending also on the actual definition of $\xi$). More specifically, using the first known terms of the $\beta$ functions, the asymptotic behavior of the length scale in the zero-temperature limit is (see, e.g., Refs. [2, 96])
@@
where $p$ is a model-dependent power. Therefore, apart from logarithmic corrections (due to $T\sim 1/\ln\xi$), the leading scaling-correction exponent is $\omega=2$. Such a behavior has been verified explicitly in the large-$N$ limit, see, e.g., Ref. [577].
We also mention that the nature of the asymptotic low-temperature behavior of the 2D RPN-1 models has been long debated, see, e.g., Refs. [250, 248, 247, 246, 245, 244]. Refs. [246, 247, 248] reported arguments that support the claim that RPN-1 and $N$-vector models belong to the same universality class, implying the irrelevance of the ${\mathbb{Z}}_{2}$ gauge symmetry. They are essentially related to the fact that $N$-vector and RPN-1 have the same asymptotic perturbative behavior. However, extensive numerical results for several variant RPN-1 models [250] show substantial differences in the nonperturbative scaling functions, which may be explained by the nonperturbative relevance of the topological ${\mathbb{Z}}_{2}$ defects characterizing the RPN-1 models, giving rise to distinct universality classes.
We finally mention that 2D systems with an Abelian O(2) global symmetry, such as the $N=2$ XY vector model and the RP1 model, are peculiar in this respect, since they can undergo a finite-temperature topological BKT transition [457, 458, 459], which separates the high-$T$ disordered phase from the low-temperature nonmagnetized phase characterized by spin waves and quasi-long range order with vanishing magnetization. At BKT transitions the correlation length $\xi$ scales as $\xi\sim\exp(c/\sqrt{T-T_{c}})$ approaching the BKT critical temperature $T_{c}$ from the high-temperature phase. The scaling behavior at BKT transitions has been extensively discussed in the literature, see, e.g., Refs. [457, 458, 459, 484, 485, 486, 487, 96, 488, 200].
@@ A.6 Finite-size scaling at first-order transitions
The behaviors emerging at first-order phase transitions are more complex than those observed at continuous transitions, see, e.g., Refs.[116, 118]. Critical phenomena at continuous phase transitions are essentially related with the presence of critical correlations, which decay as a power of the distance at the critical point, and with a diverging length scale $\xi$. When approaching the critical point, the long-distance behavior, i.e., on distances of the order of the scale $\xi$, shows universal features that only depend on few global properties of the microscopic short-distance interactions. The behavior at first-order transitions is instead more complex. Indeed, we typically observe the coexistence of different phases, with one or more phases that are ordered on all length scales, apart from some short-range local fluctuations. The presence of coexisting phases gives rise to peculiar competing phenomena in the phase-coexistence region, such as metastability, nucleation, droplet formation, coarsening, etc. [116, 578]. Moreover, the observed bulk behavior crucially depends on the nature of the boundary conditions, at variance with what happens at continuous transitions, where boundary conditions only affect some finite-size properties of the system, but are irrelevant in the thermodynamic limit. The sensitivity to the boundary conditions represents one of the main qualitative differences between continuous and first-order phase transitions [118].
@@
To distinguish continuous from first-order transitions in numerical studies of systems with limited size, one can exploit the substantial differences of their FSS behavior. A possible approach consists in considering the large-$L$ behavior of the specific heat or of the susceptibility of the order parameter, that should have a maximum that diverges as $L^{d}$[587, 588, 589, 590] In practice, one should compute the maximum of the susceptibility $\chi_{A,\rm max}(L)$ of a local operator $A_{\bm{x}}$ for each value of $L$ and then fit the results to a power law $aL^{p}$. If one obtains $p\approx d$, the transition is of first order; otherwise, the transition is continuous, $p$ being related to standard critical exponents. If $A$ is the energy density, then $\chi_{A}$ is the specific heat that diverges as $L^{\alpha/\nu}$ (if $\alpha$ is positive), while if $A$ is the magnetization, then $\chi_{A}\sim L^{2-\eta}$. This approach works nicely for strongly discontinuous transitions. In the case of weakly discontinuous transition, however, the asymptotic behavior $\chi_{A}\sim L^{d}$ may set in for values of $L$ that are much larger than those considered in the simulations. Thus, data may show an effective scaling $\chi_{A,\rm max}(L)\sim L^{p}$, with $p$ significantly smaller than $d$, effectively mimicking a continuous transition (the Potts model [591, 592] in two and three dimensions is a good example, see Refs. [587, 588, 589, 590]).
@@
An extended review of the FSS behavior at thermal and quantum first-order transitions and of its dependence on the boundary conditions can be found in Ref. [118].
@@ Appendix B Critical exponents of some O($N$)-vector universality classes
In this appendix we report the presently most accurate estimates of the critical exponents for the most common 3D $N$-vector universality classes, the Ising ($N=1$), XY ($N=2$), and Heisenberg ($N=3$) universality classes. An effective LGW model is the $\Phi^{4}$ field theory for an $N$-component real field, while a lattice representative is the standard $N$-vector model defined by the Hamiltonian (1). The estimates are reported in Tables 1, 2, and 3 for $N=1$, $N=2$, and $N=3$, respectively. A more complete list of theoretical and experimental results can be found in Ref. [96] (at least up to 2002); those reported here are the most accurate in the various approaches. We recall that the critical exponents controlling the asymptotic exponents of other observables can be obtained from $\nu$ and $\eta$ by using scaling and hyperscaling relations (see, e.g., Refs. [2, 115]). Accurate estimates of the critical exponents of the O($N$) vector universality classes for $N\geq 4$ can be found in Refs. [322, 96].
@@
Table 1: Estimates of the critical exponents of the 3D Ising universality class (characterized by a global ${\mathbb{Z}}_{2}$ symmetry). We report the correlation-length exponent $\nu$, the order-parameter exponent $\eta$, and the exponent $\omega$ associated with the leading scaling corrections. We report: SFT results obtained by resumming high-order perturbative expansions (6-loop calculations in the fixed-dimension scheme [149, 594, 595, 596, 148, 597] and in the $\varepsilon$-expansion scheme [161, 598, 599, 600]), by using functional RG approach [164], and the conformal field theory (CFT) bootstrap [159, 520] approach; results obtained in lattice models by resumming high-temperature expansions [150] and in MC simulations [156, 593, 601, 602]. We note that there is an overall agreement among the results obtained by the different approaches. Moreover, there is also a good agreement with experiments in physical systems that undergo continuous transitions in the Ising universality class (liquid-vapor systems, binary systems, uniaxial magnetic systems, Coulombic systems, etc.) see, e.g., Ref. [96] for a list of experimental results.
@@
Table 2: Estimates of the critical exponents of the 3D XY universality class (characterized by a global O(2) symmetry). We report the correlation-length exponent $\nu$, the order-parameter exponent $\eta$, and the exponent $\omega$ associated with the leading scaling corrections. We report: SFT results obtained by resumming high-order perturbative expansions [149, 161] (6-loop calculations in the fixed-dimension and in the $\varepsilon$-expansion scheme), by using functional RG approach [164] and the conformal field theory (CFT) bootstrap [165] approach; results obtained in lattice models by combining high-temperature and MC results [154] and in MC simulations [153, 163]. We also mention that experimental estimates in a microgravity environment have been obtained for the 4He superfluid transition [603, 604, 605]. See Ref. [96] for a more complete list of theoretical and experimental results.
@@
Table 3: Estimates of the critical exponents of the 3D Heisenberg universality class (characterized by a global O(3) symmetry). We report the correlation-length exponent $\nu$, the order-parameter exponent $\eta$, and the exponent $\omega$ associated with the leading scaling corrections. We report: SFT results obtained by resumming high-order perturbative expansions [149, 161] (6-loop calculations in the fixed-dimension and in the $\varepsilon$-expansion scheme), by using functional RG approach [164], and the conformal field theory (CFT) bootstrap [166] approach; results obtained in lattice models by combining high-temperature and MC results [606, 157] and in MC simulations [167]. See Refs. [96, 167] for a more complete list of theoretical and experimental results.
Appendix C Mean-field analysis of the low-temperature Higgs phases
In this appendix we present a mean-field analysis of the low-temperature Higgs phases of the LNAH model with SU($N_{c}$) gauge symmetry and $N_{f}$ degenerate scalar flavors in the fundamental representation. This model has been discussed in Sec. 9.1. For more details see Ref. [100]. This analysis allows one to infer the symmetry properties of the low-temperature Higgs phases, which depend on the parameter $v$, the number of colors $N_{c}$ and of flavors $N_{f}$. It is worth noting that this discussion applies to generic $d$-dimensional systems, therefore also to 4D space-time systems that may be relevant in the context of high-energy physics. Analogous analyses can be performed for SU($N_{c}$) theories with scalar matter in the adjoint representation [97, 289, 99], and SO($N_{c}$) gauge LNAH models [408].